OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–1, 3) and (1, 7)? y – 7 = 4(x – 1) y – 7 = 2(x – 1) y – 3 = 2(x – 1) y – 3 = 4(x + 1)

OpenStudy (waleed_imtiaz):

U know point slope form is y-y1 =m(x-x1), so here U see, We dont know the slope...... SO use the formula for slope... slope= y2-y1/x2-x1 so x1=-1 x2=1 y1=7 y2=7 So we get the slope=2 now just put the point, Any point U like But according to the options, second point is used i.e. (1,7).. U will get eh answer B y-7=2(x-1)

OpenStudy (waleed_imtiaz):

Sorry y1=3

OpenStudy (anonymous):

The easiest way to do problems like these is in 2 steps: you need the answer in y=mx+b form. To start, we'll find m which is the slope of the line using the formula: \[m=(y2-y1)/(x2-x1)\] where the points you are given are (x1, y1) = (-1,3) and (x2,y2)=(1,7) Plugging that in, you get m=2. To get it into point slope form (ie. (y-a)=m(x-b)) you just put in a point. So you would end up with the answer: y-3=2(x+1) The closest one is B as the other guy pointed out. But it should be 2(x+1) on the right side.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

My bad, it is B with the 2nd point

OpenStudy (waleed_imtiaz):

@rkfitz053 ... Here U can see.... the equation U calculated isn't in the options. So u can try the second point.... It will give U the answer B.